OpenStudy (javyb13):

WILL FAN AND MEDAL If the inverses of two functions are both functions, will the inverse of the sum or difference of the original functions also be a function? also provide 2 examples

OpenStudy (javyb13):

@AloneS @mathmate @sweetburger

OpenStudy (javyb13):

@Owlcoffee

satellite73 (satellite73):

suppose \(f(x)=2x+1\) and \(g(x)=2x-3\) both have inverses because they are lines what is \(f-g\)?

OpenStudy (javyb13):

4?

OpenStudy (javyb13):

@satellite73

OpenStudy (baayen):

Do you still need help with this one

OpenStudy (javyb13):

yes

OpenStudy (baayen):

So lets take the example provided earlier, f(x)=2x+1, and g(x)=2x+3, \[f ^{-1}(x)=1/2x -1/2 \] and \[g ^{-1}(x)=1/2x-3/2 \], So yes the inverses are a function

OpenStudy (javyb13):

i see

OpenStudy (baayen):

And now it is asking if the inverse of the sum of two functions would be a function, So we have f(x)-g(x) which would simply be f-g(x)=2x+1-2x+3, can you simplify that for me?

satellite73 (satellite73):

hold on

satellite73 (satellite73):

\[f-g=4\] in the example i gave, a number

satellite73 (satellite73):

a number (constant function) does not have an inverse

OpenStudy (baayen):

Sorry it should be f-g(x)=2x+1-2x-3

OpenStudy (baayen):

And that would be for the difference of the functions the sum of the functions would be f-g(x)=2x+1+2x+3

OpenStudy (baayen):

So what would that tell you about the sum of the functions and the difference of those given functions

OpenStudy (javyb13):

that they are both functions?

OpenStudy (baayen):

No, that would tell you that the sum of the functions is but the difference of the functions is not for the example given